v=11sin4t-5cos5t


1
Expert's answer
2021-04-29T17:13:28-0400

v=11sin4t-5cos5t


Differentiate w.r.t. t and we get-


"\dfrac{dv}{dt}=11\times 4cos4t+25sin 5t"


"\dfrac{dv}{dt}=44cos4t+25sin5t"